A vendor sells hot dogs and bags of chips. A customer buys 2 hot dogs and 2 bags of chips for $7.50. Another customer buys 5 hot dogs and 3 bags of chips for $17.25. Find the cost of each item
Answer by ewatrrr(24785)


Hi
2x + 2y = 7.50
0r
x + y = 3.75 x(-3)
5x + 3y = 17.25
-3x - 3y = -11.25 x(-3)
5x + 3y = 17.25
2x = 6
x = $3.00, y = 0.75
CHECKING our answer***Using 5x + 3y = 17.25
$15.00 + $2.25 = $17.25
Wish You the Best in your Studies.